Your firm has an average collection period of 22 days. Current practice is to factor all receivables immediately at a discount of 1.2 percent. Assume that default is extremely unlikely.

What is the effective cost of borrowing?
Answer #1

Interest rate = discount rate/(1-discount rate)

=1.20%/(1-0.012)= 1.20%/0.988

=1.2145749% or 1.21%

N= 365/average condition period = 365/22

N=16.59

Effective cost of borrowing

EAR=(1+I)^n-1

= (1+0.0121)^12.59-1

= 1.16349 -1= 16.35%